I have 2 more simultaneous questions that I need help with please. Very confused. Thank you
5x+y=19
2x-y=2
and
6x+7=20
3x+y=11

Answer by mathman409(67)

5x+y=19
2x-y=2
add the equations to cancel 1 variable out, then figure out the other variable.
5x+y=19
+
2x-y=2
____________
7x=21
x=3
get the other variable using any equation, but this time replace x with the value it has, in this case, 3
15+y=19
y=19-15
y=4
CHECK USING THE OTHER EQUATION (this is a very important step if you want to be good at algebra)(assume your answers are right and use the other equation. replace variables with their assumed correct value. if you have the same number on both sides of the equal sign, than you're correct. if you don't...well, than-- you'll just have to start over again)
2(3)-4=2
6-4=2
2=2
yay we did it!